2(x-1)/3=1+3x-(6/4)
Answer by kuhal(1)


= 2(x-1)/3=1+3x-(3/2)
by mult. (3)(2) we get :
4(x-1)=2+6x-3
4x-4=6x-1
4x-6x=4-1
-2x=3
x=-3/2